# Quantum Vacuum energy as the origin of Gravity

### v2

@@ I Introduction
Gravity remains the least understood of the fundamental forces, in spite of the discoveries of General Relativity (GR) and Quantum Mechanics (QM) over 100 years ago, this year being recognized as the 100-th anniversary of a complete theory of QM due to Schrödinger and Heisenberg, although the birth of QM goes back to Planck’s black body studies at finite temperature. Classical singularities at the center of black holes and at the time of the Big Bang have not yet been understood and resolved in classical GR. Many physicists expect that such singularities will be resolved by a theory of quantum gravity, but it’s fair to say that a complete theory of quantum gravity has not been forthcoming, although string theory offers some promise. For these reasons it is worthwhile to pursue new ideas on the origin or emergence of Gravity itself, especially if such ideas suggest that perhaps Gravity doesn’t need to be quantized in the first place. Such a theory should be able to explain the value of Newton’s constant $G_{N}$ itself.
@@
In this article we explore the question “Can Quantum vacuum energy be the origin of Gravity?” [^2] This article was initially intended simply as a theoretical exercise in the form of an essay to address this question, but we found it can lead to a fruitful new perspective on Gravity more generally, and it is not without some predictions, if the basic ideas continue to hold. This perspective is based on the idea that perhaps we can learn more about the foundations of Gravity by considering the cosmological far future, in contrast to the far past where much less is understood, in particular the origin of the Big Bang and possibly inflation. The late Universe is dominated by so-called Dark Energy, thus one should first attempt to understand the full implications of this in a Universe where Dark Energy is its sole component. By assumption, we will equate Dark Energy with the quantum vacuum energy density $\rho_{\rm vac}$} computed in quantum field theory (QFT) in flat local Minkowski space, defined as
@@
with the convention $\{\eta_{\mu\nu}\}={\rm diag}\{1,-1,-1,-1\}$ where the Vacuum $|0\rangle$ is the ground state for all quantized fields, excluding gravity. The gravitational implications of the quantum field theoretic properties of $\rho_{\rm vac}$} will be then explored.
@@
The first observations of Gravity were of course at the surface of the Earth which is manifested as a constant acceleration $g$ for all matter regardless of mass or specific material. Newton later understood $g$ to be “environmental” rather than fundamental, namely $g=G_{N}M_{E}/R_{E}^{2}$, where $M_{E},R_{E}$ are the mass and radius of the Earth. In Newton’s universal theory of Gravity, $G_{N}$ is considered a new fundamental constant of Nature, and this view was adopted by Einstein in his formulation of General Relativity. [^3] Throughout this paper, $G_{N}=6.67\cdot 10^{-11}{\rm m^{3}}/{\rm kg\,s^{2}}$ is Newton’s constant as measured today. We keep the fundamental constants $\hbar$ and $c$ explicit in order to make certain conceptual points and to provide some numbers with units. “Natural” units based on the Planck length $\ell_{p}$ can be deceptive and are not so natural for this article. In contrast, below we will argue that $G_{N}$ is also environmental, namely
@@
Based on the above remarks, what will be central to our exercise is the quantity $\rho_{\rm vac}$}. For purposes of illustration, consider a Universe which only consists of a single harmonic oscillator with ground state energy $E_{0}=\hbar\omega/2$, in addition to an imagined observer. Since classical mechanics in unaffected by shifts of the potential energy by an arbitrary constant, $E_{0}$ is not measurable in classical physics. If all that exists is the harmonic oscillator in empty space, then $E_{0}$ also cannot be measured in a single experiment based on QM. There are two known ways to measure $E_{0}$. First, one can couple the harmonic oscillator to a generic heat bath at temperature $T$. Then one can in principle infer $E_{0}$ from the high temperature behavior of the partition function or its logarithm which is the free energy. It’s important that this is based on the free energy and not its derivatives, and thus it is not invariant under shifts of the potential energy. The second manner to measure $E_{0}$ is to couple the oscillator to Gravity, since gravity originates from all the matter-energy in the Universe and is sensitive to its ground state energy. The fact that the ground state energy can be measured either by coupling to a heat bath or gravity suggests parallels between thermodynamics and gravity that have already been recognized. Deceptively simple as it may appear, to study QFT at finite temperature with periodic boundary conditions, one necessarily needs to study quantum fields in curved Euclidean space. In 1 spatial dimension the Euclidean space is a torus, and the problem of conformal theory on a torus is a completely solved problem. Like Gravity, the Vacuum $|0\rangle$ is omnipresent. The Vacuum “knows” the physical properties, such as mass, quantum numbers, and interactions of any particle excitations above it, and is perhaps the only entity that exists everywhere in the Universe and for all times. Everything else is some quantum excited state, including people, and typically unstable. Nature does not abhor a Vacuum, rather it depends on it. [^5] To paraphrase with translation a recognizable philosopher, If you look deeply enough into the Void, it eventually looks back at you. So it is reasonable to suppose that the Vacuum, in particular its ground state energy density $\rho_{\rm vac}$}, is at the origin of gravity. In other words, if the only way to measure $E_{0}$ is with gravity, then reversing the argument, perhaps quantum vacuum energy is the origin of Gravity. For the same reason that one does not necessarily have to quantize the heat bath to study finite temperature QFT, perhaps one does not have to quantize Gravity, however this deep issue will be left as an open question. This does not preclude the existence of classical gravitational waves that have been measured, as they can be viewed as classical perturbations of the Vacuum as an elastic medium. This article explores this possibility as a theoretical academic exercise to see where it leads.
This article assumes that $\rho_{\rm vac}$} is well-defined in Minkowski space, namely it is as well defined as the ground state energy of a harmonic oscillator. Early considerations of $\rho_{\rm vac}$} based on viewing a free quantized field as a collection of harmonic oscillators with energies $\hbar\omega_{\bf{k}}$, with $\omega_{\bf{k}}=\sqrt{{\bf{k}}^{2}+m^{2}}$ led to the introduction of a cut-off in $|{\bf{k}}|$. If this cut-off is taken to be based on the Planck length $\ell_{p}\equiv\sqrt{\hbar G_{N}/c^{3}}$ then $\rho_{\rm vac}$} is much too high by even 100 orders of magnitude and this is perhaps the first formulation of the “cosmological constant problem” Weinberg . This naive computation of $\rho_{\rm vac}$} depends on the cut-off and goes to infinity as the cut-off goes to infinity. This computation is essentially meaningless, as is the often quoted “worse prediction in physics by $10^{120}$ orders of magnitude”, simply because it is not a real prediction. In quantum field theory one is accustomed to dealing with ultra-violet (UV) singularities. In recent work ALrhovac1 ; ALrhovac2 we formulated a proper computation of $\rho_{\rm vac}$} which renders it finite and well-defined. Since this will be central to subsequent arguments we list its main properties.
(i)   For 2 dimensional integrable theories $\rho_{\rm vac}$} is exactly computable from the Thermodynamic Bethe Ansatz, and also from the S-matrix bootstrap, and takes the form ${\rho_{\rm vac}}=m_{1}^{2}/2\mathfrak{g}$ where $m_{1}$ is the lightest mass particle. In the works ALrhovac1 ; ALrhovac2 this formula was extended to 4 spacetime dimensions:
where $m_{\rm z}$ is the physical mass of the lightest particle, which we will refer to as the “zeron”, where zero signifies the Vacuum, and $\mathfrak{g}$ is a dimensionless interaction coupling constant. Although one does not have the powerful simplifications of integrability in 4 dimensions, this formula was well motivated in ALrhovac1 ; ALrhovac2 and should not be viewed as an ad hoc assumption, although unlike in 2 dimensions it hasn’t yet been fully rigorously proven since the $1/\mathfrak{g}$ dependence indicates it is highly non-perturbative. Clearly the above expression is purely quantum mechanical as can be seen from its $\hbar$ dependence. The $1/\mathfrak{g}$ dependence can be anticipated, and is in fact necessary, since $\rho_{\rm vac}$} should diverge as $\mathfrak{g}\to 0$ due to an unavoidable logarithmic divergence for free QFT in the UV when the spatial dimension $d$ is evenALrhovac1 . [^6] This is essentially another, but refined, version of the “Cosmological Constant Problem”. The precise identification of the zeron as a real particle will not be necessary for our arguments, thus this identification will be left as an open question. However it was argued in ALrhovac2 that $m_{\rm z}$ is fixed in the deep UV above the electro-weak spontaneous symmetry breaking scale, and that Majorana neutrinos are viable candidates (see (8)). [^7] For experimental constraints on neutrino masses see for instance NeutrinoMasses .
[…]
for some constant coefficients $c_{\rm uv},c_{4}$. [^8] In 2 spacetime dimensions $c_{\rm uv}$ is the effective Virasoro central charge. Then ${\rho_{\rm vac}}=-c_{4}\pi^{2}m_{\rm z}^{4}/90$. This was used in ALrhovac1 to estimate $\rho_{\rm vac}$} for QCD based on finite temperature lattice results.
(iii)   $\rho_{\rm vac}$} can be independently computed from the form-factor bootstrap that relates the two particle form-factor to the zero particle one, which is a 1-point vacuum expectation value of the field in question ALrhovac2 . This is the reason for $m_{\rm z}$ to be the lightest mass particle, since in principle the S-matrix and form factors for other particles can be obtained by bootstrapping the lightest mass particle. Applying this to $\langle 0|T_{\mu\nu}|0\rangle$, this leads to a determination of the coupling $\mathfrak{g}$. As usual, defining $\cal T$} from the S-matrix $S=1+i{\cal T}$, then
where $s=(p_{1}+p_{2})^{2}$ is the Mandelstam variable for 2-particle scattering.
In two spacetime dimensions, the two above definitions of $\rho_{\rm vac}$} were shown to agree exactly for a variety of integrable quantum field theories ALrhovac1 ; ALrhovac2 . As previously mentioned, in one spatial dimension it takes the form ${\rho_{\rm vac}}=c^{3}m_{1}^{2}/2\hbar\mathfrak{g}$ where $m_{1}$ is the lightest particle and $\mathfrak{g}$ can be explicitly computed. Various consistency checks on this formula were made, including that it must vanish for supersymmetric theories and also theories with a fractional supersymmetries where the hamiltonian is in the center of the universal enveloping algebra of the conserved charges. (See ALrhovac1 and references therein.) The coupling $\mathfrak{g}$ can be positive or negative, and $\rho_{\rm vac}$} can even oscillate in sign for a given model, for instance for the sine-Gordon model where it oscillates from positive to negative, where it is zero at the fractional supersymmetric points. In 4 spacetime dimensions it was not rigorously shown that equations (4), (5) lead to the same formula for $\rho_{\rm vac}$}, thus (3) should still be viewed as well-motivated but still conjectural. For the remainder of the article we simply assume the well-motivated formula (3).
Some independent support for the result (3) comes from Swampland ideas Festina1 ; Festina2 . By studying a charged particle of mass $m$ in the presence of a black hole, it was argued that
${\rho_{\rm vac}}<\dfrac{m^{4}}{2e^{2}}$ (6)
where here $\mathfrak{g}$ is identified with the quantum electrodynamic (QED) fine structure constant $\alpha=e^{2}/\hbar c$. Actual QED will not play a role in this article since we assume the zeron is electrically neutral, however, as we will see, it will be instructive to compare couplings and renormalization group parameters with QED, just to check if they are reasonable. Our result (3) is stronger than (6) since it is not an inequality, but it is still consistent with (6).
As previously stated, in this article we do not propose what the hypothetical zeron particle actually is, nor even that it is a detectable particle, however we can constrain it’s mass $m_{\rm z}$ from the cosmologically measured value of $\rho_{\rm vac}$}. Taking it to be ${\rho_{\rm vac}}=\Omega_{\Lambda}\rho_{\rm crit}$, where $\Omega_{\Lambda}\approx 0.68$ one finds Planck
${\rho_{\rm vac}}=5.2\times 10^{-10}\,\dfrac{{\rm Joule}}{{\rm m}^{3}}.$ (7)
Using this value, then (3) implies
$m_{\rm z}=0.0024\,\mathfrak{g}^{1/4}\,{\rm eV}.$ (8)
For $\mathfrak{g}={\cal O}(1)$, it is interesting to note that $m_{\rm z}$ is on the order of proposed neutrino masses NeutrinoMasses . Below we will propose additional constraints on the coupling $\mathfrak{g}$.
@@
It is instructive to compare the perspective on Gravity developed in this article with some proposed theories of quantum gravity, since this clearly delineates some important distinctions. We take “quantum gravity” to mean a theory where the space-time metric $g_{\mu\nu}(x)$ itself is quantized which implies the existence of spin 2 gravitons. Although the definition of the Planck length $\ell_{p}$, defined from $G_{N}$ as $G_{N}=c^{3}\,\ell_{p}^{2}/\hbar$, is a priori just based on dimensional analysis, one expects that $\ell_{p}$ is a relevant scale when considering graviton-graviton scattering. The latter does not play a role in this article, which makes one question whether $\ell_{p}$ is a fundamental constant, however this issue will not be completely settled in the present cosmological context. [^9] Based on this article, units based on $\hbar$, $c$ and $m_{\rm z}$ are more natural. A significant result of string theory is that it provides a consistent prescription for the computation of graviton scattering that is finite StringTheory . Thus gravity “emerges” from string theory, in the sense that Newton’s constant $G_{N}$ is determined by matching graviton scattering with low energy gravity. This leads to a formula for $G_{N}$:

### v3

@@ I Introduction
Gravity remains the least understood of the fundamental forces, in spite of the discoveries of General Relativity (GR) and Quantum Mechanics (QM) over 100 years ago, this year being recognized as the 100-th anniversary of a complete theory of QM due to Schrödinger and Heisenberg, although the birth of QM goes back to Planck’s black body studies at finite temperature. Classical singularities at the center of black holes and at the time of the so-called Big Bang have not yet been understood and resolved in classical GR. Many physicists expect that such singularities will eventually be resolved by a theory of quantum gravity, but it’s fair to say that a complete theory of quantum gravity that can address these fundamental issues has not been forthcoming. String theory still offers some promise and is the most well-motivated since it provides a consistent scattering theory of gravitons and more. Nevertheless, for these reasons it is worthwhile to pursue new ideas on the origin or emergence of Gravity itself, especially if such ideas hint that perhaps Gravity doesn’t need to be quantized in the first place, or at the very least is not yet needed to explain currently observable phenomena. Such a theory should be able to explain the value of Newton’s constant $G_{N}$ itself.
Like Gravity, the Vacuum $|0\rangle$ is omnipresent. The Vacuum “knows” the physical properties, such as mass, quantum numbers, and interactions of any particle excitations created above it, and is perhaps the only entity that exists everywhere in the Universe and for all times. The existence of the Quantum Vacuum guarantees that the allowed particle excitations above it, such as electron/positron pairs, are identical throughout the Universe, because they already exist virtually in both a physical and mathematical sense. Everything else we observe are quantum excited states of the Vacuum above the ground state it represents, including people, and are typically unstable. Nature does not abhor a Vacuum as Aristotle first stated, rather it depends on it. [^2] To paraphrase with translation a recognizable philosopher, If you look deeply enough into the Void, It eventually looks back at you. So it is reasonable to suppose that the Vacuum, in particular its ground state energy density $\rho_{\rm vac}$}, is at the origin of Gravity. If the only way to measure the ground state energy $E_{0}$ is with Gravity, then reversing the argument, perhaps quantum vacuum energy is the origin of Gravity. [^3] In a certain sense, the old aether is back as the Vacuum itself. Deceptively simple as it may appear, to study quantum field theory (QFT) at finite temperature $T$ with periodic spatial boundary conditions, one necessarily needs to study quantum fields in curved euclidean space where the euclidean time is compactified to a circle with circumference $\hbar c/k_{B}T$. In one spatial dimension the euclidean spacetime is a torus, and the problem of $d=1$ conformal QFT on a torus is completely solved, even though it is a non-trivial problem involving modular invariance, etc. From this point of view, temperature is a fundamental link between geometry and QFT. For the same reason that one does not necessarily have to quantize the heat bath to study finite temperature QFT, perhaps one does not have to quantize Gravity. If the Vacuum and its energy density is the origin of Gravity itself, then since the Vacuum is intrinsically Quantum, one ponders whether quantizing gravity is essentially redundant. Henceforth, we set aside the deep question as to whether Gravity ultimately needs to be quantized since, as we will see, a resolution of the singularity at the origin of the so-called Big Bang can be understood without it. [^4] We take “quantum gravity” to mean quantization of the metric gravitational field itself, which necessarily implies spin-2 massless gravitons. Other arguments that attempt to prove that gravity must be quantized are not based on gravitons, but rather on the principles of superposition and unitarity in QM. These arguments attempt to prove that a purely classical gravitational field leads to inconsistencies with the rest of the quantized world. In a somewhat fanciful thought experiment put forward by Feynman Feynman , imagine a massive ball that becomes entangled with an electron, where the electron is in a superposition of two states with different paths, thus the ball becomes a superposition of states with different locations. This would imply the gravitational field of the ball is in a superposition, and this could in principle be detected by interacting gravitationally with a second ball. These arguments rely on unitarity of QM, and whether the Universe is an open or closed system is debatable. For instance, in the search for a de Sitter/CFT correspondence, it has been proposed that the CFT should be non-unitary Strominger . Recently a potential loop-hole in Feynman’s thought experiment has been presented Oppenheim . This does not preclude the existence of classical gravitational waves that have been measured, as they can be viewed as classical perturbations of the Vacuum as an elastic medium.
For purposes of illustration, consider a universe which only consists of a single harmonic oscillator in its ground state with energy $E_{0}=\hbar\omega_{0}/2$, in addition to an imagined observer. Since classical mechanics is unaffected by shifts of the potential energy by an arbitrary constant, $E_{0}$ is not measurable in classical physics. If all that exists is the harmonic oscillator in empty space, then $E_{0}$ also cannot be measured in a single measurement based on QM, but can in principle be measured if coupled for instance to photons with repeated measurements. There are two known general ways to measure $E_{0}$. First, one can couple the harmonic oscillator to a generic heat bath at temperature $T$. Then one can in principle infer $E_{0}$ from the high temperature behavior of the partition function or its logarithm which is the free energy. It’s important that this is based on the free energy and not its derivatives, and thus it is not invariant under shifts of the potential energy, and this avoids some fine-tuning issues. The second manner to measure $E_{0}$ is to couple the oscillator to Gravity, since gravity originates from all the matter-energy in the Universe and is sensitive to its ground state energy. The fact that the ground state energy can be measured either by coupling to a heat bath or gravity suggests parallels between thermodynamics and gravity that have already been recognized, in particular for the entropy of black holes. If our imagined observer of the single harmonic oscillator thinks about natural units, they would be based on $\hbar$, $c$, and $E_{0}$ which are complete, as they lead to a fundamental length, time and mass.
@@
Based on the above remarks, the initial perspective on Gravity taken in this article is based on the idea that perhaps we can learn more about the foundations of Gravity by considering the cosmological far future, in contrast to the far past where much less is understood, in particular the origin of the Big Bang and possibly inflation. The late Universe is dominated by so-called Dark Energy, thus one should first attempt to understand the full implications of this in a Universe where Dark Energy is its sole component. Below we will refer to such a hypothetical universe as the dark-universe, and it should be viewed as a skeleton of a model of the Universe where there are no excitations over its ground state. By assumption, we will equate Dark Energy with the quantum vacuum energy density $\rho_{\rm vac}$} computed in quantum field theory (QFT) in flat local Minkowski space, defined as
@@
with the convention $\{\eta_{\mu\nu}\}={\rm diag}\{1,-1,-1,-1\}$ where the Vacuum $|0\rangle$ is the ground state for all quantized fields, excluding gravity. The gravitational implications of the quantum field theoretic properties of $\rho_{\rm vac}$} will be then explored. Thus this article essentially explores the question “Can Quantum vacuum energy be the origin of Gravity?” The ideas presented here are most closely related to Andrei Sakharov’s idea on “induced gravity” Sakharov ; Visser , however they differ in many important details. [^5] The parallels with Sakharov’s induced gravity ideas were only pointed out to us after the first version of this article was made available on arXiv. Sakharov’s original article is extremely short with very little equations and is difficult to penetrate. A nice explanation can be found in MisnerThornWheeler page 426, where it is explained how Newton’s constant $G_{N}$ is viewed as a kind of elastic modulus. For instance our model does not require a UV cut-off in momentum space as Sakharov’s theory does, since this UV divergence is controlled in the QFT through renormalization, and this renders $\rho_{\rm vac}$} well-defined both physically and mathematically. This article can thus be viewed as applying modern ideas of the renormalization group to Sakharov’s idea to deal with the cut-off, which were not formulated until after Sakhavov’s work by Wilson KWilson , and the resulting renormalization group flow for $\rho_{\rm vac}$}. The theoretical framework is still semi-classical gravity where $T_{\mu\nu}$ for Dark Energy is identified with its vacuum expectation value. We are unaware of applications of Sakharov’s ideas to the Hubble tension in the literature, nor its prediction of the minimal scale factor $a_{\rm min}$ described below.
The main assumptions in this article are rather conservative and minimal, and it’s useful to itemize them here in order to understand how they fit into the existing literature.
Assumption 1.    Dark Energy, or equivalently the Cosmological Constant, is equated with quantum vacuum energy density in flat Minkowski space, namely the vacuum expectation value of the energy-momentum tensor (1). This quantity is finite and well-defined in any QFT and studied in some detail recently for 4 dimensional spacetime ALrhovac1 ; ALrhovac2 . There we obtained the formula
where $m_{\rm z}$ is the physical (renormalized) mass of the lightest massive particle, and $\mathfrak{g}$ is a dimensionless interaction coupling constant, or a function thereof if higher order corrections in perturbation theory are included. The above quantity refers to the vacuum energy density for all known particles, and the fact that the lightest mass particle $m_{\rm z}$ appears is based on bootstrap principles. This thus assumes there is a single particle mass scale that determines all other particle masses. This is actually the situation in the Standard Model of particle physics where all masses are thought to arise from the single scale through the Higgs mechanism. This assumption is even more correct if there is a Grand Unified Theory of particle physics. If the Vacuum and its energy density $\rho_{\rm vac}$} were based on two or more completely independent and decoupled QFT’s with different fundamental mass scales, then the formula (2) doesn’t obviously apply since it becomes undetermined what $m_{\rm z}$ actually is. Thus we have to assume all particles can interact with each other, otherwise one is dealing with separate, distinct universes. The formula (2) is not ad hoc, and was well-motivated in ALrhovac1 ; ALrhovac2 . Since this formula is central to our study, its underpinnings are reviewed and explained below in Section IVA. It is not necessary for our purposes at this preliminary stage to definitively identify the $m_{\rm z}$-particle as a known particle, however we can constrain its mass, equation (5), and it’s coupling constants based on Hubble tension data, and we will have more to say about its properties below.
Assumption 2.    The coupling $\mathfrak{g}$ is a marginally irrelevant coupling with 1-loop renormalization group (RG) beta function:
$\mu\partial_{\mu}\mathfrak{g}=\dfrac{b}{2\pi}\mathfrak{g}^{2}$ (3)
for some positive constant coefficient $b$, where increasing $\mu$ corresponds to a flow to higher energies. Integrating the RG flow leads to the parameter $\widehat{b}=b\,\mathfrak{g}_{0}/2\pi$ where $\mathfrak{g}_{0}$ is the value of $\mathfrak{g}$ at the relevant energy scale today. The parameter $\widehat{b}$ is what modifies the Friedmann equations in the way presented below. Based on the Hubble tension we will constrain $\widehat{b}\approx 0.02$ which is quite small, and justifies a 1-loop approximation.
Assumption 3.    There exists an energy scale of the Universe corresponding to a time dependent temperature $T(t)$, where as usual
$T(t)=\dfrac{T_{0}}{a(t)}=T_{0}\,(1+z(t))$ (4)
where $a(t)$ is the scale factor and $z$ the redshift, and $T_{0}$ is the temperature today. As we will see, de Sitter space has a natural temperature (29). We assume $T_{0}$ is the temperature of the Cosmic Microwave Background (CMB) today, $T_{0}\approx 2.7$K, however some predictions are not so sensitive to the exact value of $T_{0}$ and it could in principle be the temperature based on neutrinos which is not so different. We mainly focus on the $\Lambda{\rm CDM}$ model, but we will assume that $T(t)$ is meaningful at energy scales above that of the CMB, namely $z>1100$, at least up to some limit $T_{\rm max}\approx T_{0}(1+z_{\rm max})$ where $z_{\rm max}\sim e^{1/\widehat{b}}$. (See below for explanations of $z_{\rm max}$.)
Assumption 4.    Matter and radiation are treated as particle excitations of the Vacuum above its ground state, i.e. the Vacuum, with energy density $\rho_{\rm vac}$}. We assume $\rho_{\rm vac}$} is the vacuum energy density for all quantum fields, and all of the multiple interactions and parameters of the Standard Model of particle physics are implicit in $\rho_{\rm vac}$}. This allows for various deviations of (4) at specific times where various phase transitions can occur, such as the electro-weak phase transition and QCD phase transitions. We incorporate matter and radiation after first understanding a universe composed of only vacuum energy, which we will refer to as the dark-universe, and then subsequently add matter and radiation consistent with local energy-momentum conservation $\nabla^{\mu}T_{\mu\nu}=0$, which leads to the usual Friedmann equations with all forms of energy included. As already stated, the dark-universe  should be viewed as a skeleton for a complete universe which includes quantum matter and radiation.
Assumption 5.    We don’t assume gravity is quantized, in part simply because we don’t need it yet to resolve the curvature singularity of the Big Bang, as we will show.
Under these assumptions all of the QM involved is contained in the QFT computation of $\rho_{\rm vac}$} and it’s renormalization group properties in Minkowski space. We are thus implicitly working in the framework of semi-classical gravity, and simply work out the implications of these minimal, well-founded assumptions, to see where it leads us. The parameters introduced are $m_{\rm z}$, which sets the scale of the cosmological constant, $\widehat{b}$, and the temperature $T$. The Hubble constant $H_{0}$ is just an initial condition for the Friedmann differential equations at the present time $t_{0}$, as is $T_{0}$. Taking $\rho_{\rm vac}$} to be its currently measured value in cosmology, namely ${\rho_{\rm vac}}=\Omega_{\Lambda}\rho_{\rm crit}$, where $\Omega_{\Lambda}\approx 0.68$ one finds Planck
${\rho_{\rm vac}}\approx 5.2\times 10^{-10}\,\dfrac{{\rm Joule}}{{\rm m}^{3}},~~~~~\Longrightarrow~~m_{\rm z}\approx 0.0024\,\mathfrak{g}^{1/4}\,{\rm eV}.$ (5)
For $\mathfrak{g}={\cal O}(1)$, it is interesting to note that $m_{\rm z}$ is on the order of proposed neutrino masses NeutrinoMasses , however as previously stated, the precise identification of the particle with mass $m_{\rm z}$ will not be necessary for our purposes, and is thus left as an open question. Below we will propose additional constraints on the coupling $\mathfrak{g}$. From these parameters alone one can develop a pseudo-realistic cosmos which can serve as a skeleton for our own Universe. All of the “microscopic” details of the evolution of the Universe obviously depend on the parameters, interactions, and all other bells and whistles of the complete QFT in Minkowski space of the Standard Model of particle physics. As we will see, within such a minimal framework one can at least address some of the most profound open problems of modern cosmology: the origin of the Cosmological Constant, the growing Hubble tension, the resolution of $a(t)=0$ singularities associated with the so-called Big Bang, what actually occurred before the Big Bang, why the Universe is flat, and whether current models of inflation play a necessary role or what replaces it. This framework may not provide final answers to all of these deep questions, but at the very least it is a novel approach to attempting to address them, and not without some new testable predictions that we will present below.
@@
Newton’s constant $G_{N}$ is not on the list of basic parameters $m_{\rm z},\widehat{b}$, and $T$ we just presented in the above Assumptions. In our proposed framework $G_{N}$ it is not a fundamental constant of Nature, and consequently nor is the Planck length $\ell_{p}=\sqrt{\hbar G_{N}/c^{3}}$. From this perspective, natural units should be based on $c,\hbar,{\rm and}~m_{\rm z}$ rather than $c,\hbar,{\rm and}~G_{N}$. [^6] From these, one can define a fundamental length $\ell_{\ell}=2\pi\hbar/m_{\rm z}c$. Based on the formula (5), if $\mathfrak{g}=1$, then $\ell_{\ell}=5.2$ micrometers, which is huge compared to the Planck length $\ell_{p}$. What is more fundamental are mass/energy units. Again for $\mathfrak{g}=1$, $m_{\rm z}\approx 0.0024\,{\rm eV}/c^{2}\approx 3.6\times 10^{-36}$   grams, which should be compared with the Planck mass $m_{p}=\sqrt{\hbar c/G_{N}}\approx 2\times 10^{-5}$ grams. Based on our current understanding of fundamental particle physics, such as neutrino masses, a fundamental mass of about $0.002\,{\rm eV}/c^{2}$ appears to us more natural than $m_{p}$ which is about the mass of a large amoeba or very fine grain of sand. If $G_{N}$ is not a fundamental constant, then what determines its measured value? The first observations of Gravity were of course at the surface of the Earth which is manifested as a constant acceleration $g$ for all matter regardless of mass or specific material. Newton later understood $g$ to be “environmental” rather than fundamental, namely $g=G_{N}M_{E}/R_{E}^{2}$, where $M_{E},R_{E}$ are the mass and radius of the Earth. In Newton’s universal theory of Gravity, $G_{N}$ is considered a new fundamental constant of Nature, and this view was adopted by Einstein in his formulation of General Relativity. [^7] Throughout this paper, $G_{N}=6.67\cdot 10^{-11}{\rm m^{3}}/{\rm kg\,s^{2}}$ is Newton’s constant as measured today. We keep the fundamental constants $\hbar$ and $c$ explicit in order to make certain conceptual points and to provide some numbers with units. “Natural” units based on the Planck length $\ell_{p}$ are not so natural for this article. In contrast, below we will argue that $G_{N}$ is also environmental, namely
@@
It is instructive to compare the perspective on Gravity developed in this article with some proposed theories of quantum gravity, since this clearly delineates some important distinctions. We take “quantum gravity” to mean a theory where the space-time metric $g_{\mu\nu}(x)$ itself is quantized which implies the existence of massless spin 2 gravitons. Based on this delineation, quantum gravity plays no role in the present article, and in spite of this our model avoids the usual curvature singularity when the scale fact $a(t)=0$. Although the definition of the Planck length $\ell_{p}$, defined from $G_{N}$ as $G_{N}=c^{3}\,\ell_{p}^{2}/\hbar$, is a priori just based on dimensional analysis, one expects that $\ell_{p}$ is a relevant scale when considering graviton-graviton scattering. A significant result of string theory, which continues to be its main motivation, is that it provides a consistent prescription for the computation of graviton scattering that is finite StringTheory . Thus gravity “emerges” from string theory, in the sense that Newton’s constant $G_{N}$ is determined by matching graviton scattering with low energy gravity. This leads to a formula for $G_{N}$ which is completely determined by physics at UV energy scales: